Scheduling Meeting Minutes: 11 Feb 2008

Attending: Christoph Montag, Mike Leitch, Bill Christie, Lijuan Ruan, Fulvia Pilat, Paul Sampson, Phil Pile, Yousef Makdisi, Anatoli Zelenski, Nick Franco, Mike Sivertz.

Performance:  Last night there were low intensity with high background rates.  This was due to collimator settings from d-Au that were mistakenly loaded.  Today Angelika loaded the correct settings and rates were good.  PHENIX would like to put off declaring physics until we reach 111 bunches and 1.4e11 intensity.  With the rotator ramp Angelika will have to find new collimator settings. 

Maintenance Day:   Could start with turning off the polarized source heaters at midnight, or as soon after midnight that the store goes up.  RHIC opens at 10:00-14:00 for JET work.

AGS C-20 polarimeter work: LOTO by CAS at 6:00.  AGS work from 8:00-14:00.

Siemans maintenance: 10:00–14:00

Spray pump motor:

PP2PP: behind JET

STAR:  sweep after 3 hours

PHENIX: sweep after 3 hours, entry to the North Muon Magnet area

C-20: Reset GPIB interface for HV Bias supply.  Limit switch movement. 

APEX:  The Polarimeter studies have been pulled from APEX.  The schedule is planned to run on Thursday 5:00–17:00.


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020 (leave message)